ROBOT SERVANTS

• When will we have household robots? I’m very skeptical about present robotics.

• What will they be like? As in science fiction?

• What should they be like?

• Will they have emotions? Free will?

• What’s the biggest danger? Their use in human conflicts.

A NEW AGE OF LUXURY—LIKE THE VICTORIAN RICH • My household robot will replace my housekeeping in an hour a day. What about the other 23 hours?

• Super-cleanliness and order.

• It knows where everything is?

• The teenage girl’s ladies maid.

• The porter. No more than six robots allowed. •